OpenStudy (anonymous):

32/100 in lowest terms rand a farction

OpenStudy (saifoo.khan):

8/25

OpenStudy (anonymous):

what saif said.

OpenStudy (anonymous):

lol

OpenStudy (karatechopper):

8/25 my comp is not working and i am writing prepostions with another hand and writing give me a break!

OpenStudy (anonymous):

32/100 = 16/50=8/25